### v2

@@ 3.5 Ablation Study
The IPS tokens significantly improve the performance, and the PaE module is indispensable for the IPS method to achieve SOTA results (Tab. 4). Using IPS tokens alone represents initialization of the tokens rather than being generated from the PaE. Thus, removing the PaE provides an extremely lightweight option.
Table 4: Ablation study for pattern shifting with 16 points (mDice scores). Mask tokens refer to unfreezing the mask tokens in SAM’s mask decoder. The optimal setting is in bold. ${}^{a}$ Number of Trainable Parameters.
[…]
@@ 4 Conclusions
In this paper, we propose a novel adaptive prompt generation module, Auto-Prompting Module (APM), which improves the non-fine-tuned performance of SAM in the target domain by generating optimal Euclidean prompts. In addition, we propose Incremental Pattern Shifting (IPS), which enables non-fine-tuned pattern shifting to improve SAM’s performance in unfamiliar domains, achieving SOTA and competitive results. Furthermore, we couple IPS with APM to propose the ProMISe framework, which realizes end-to-end pattern shifting. Benefiting from IPS, increasing prompt point number results in significant performance gains for SAM, which may be further extended to include scrawl, sketch or coarse prompts. More importantly, our results prove that a fine-tuning-based approach is not necessarily optimal for utilizing a foundation model like SAM.
{credits}
4.0.1 Acknowledgements
This work was supported by Natural Science Foundation of China under Grant 62271465 and Open Fund Project of Guangdong Academy of Medical Sciences, China (No. YKY-KF202206).

### v3

@@ 5.6 Ablation Study
Apart from the differences in effects brought by the individual and combined application of IPS and APM discussed in Sect. 5.2 - 5.4, we also find that modifications to mask tokens can significantly enhance the pattern-shifting ability of the SAM-based model. As demonstrated in Tab. 6, adding IPS tokens to mask tokens significantly improve the performance, and the PaE module is indispensable for the IPS method to achieve SOTA results. Using IPS tokens alone represents initialization of the tokens rather than being generated from the PaE. Thus, removing the PaE provides an extremely lightweight option.
Table 6: Ablation study for pattern shifting with 16 points (mDice scores). Mask tokens refer to unfreezing the mask tokens in SAM’s mask decoder. The optimal setting is in bold. a Number of Trainable Parameters.
Mask tokens IPS tokens PaE Frozen SAM TPa Kvasir EndoScene ColonDB ETIS
@@ 6 Conclusions
In this paper, we propose a novel adaptive prompt generation module, Auto-Prompting Module (APM), which improves the transfer-free performance of SAM in the target domain by generating optimal Euclidean prompts. In addition, we propose Incremental Pattern Shifting (IPS), which enables non-fine-tuned pattern shifting to improve SAM’s performance in unfamiliar domains, achieving SOTA and competitive results. Furthermore, we couple IPS with APM to propose the ProMISe framework, which can realizes end-to-end pattern shifting to improve training efficiency and stability. We conducted experiments in endoscopic and dermoscopic benchmark datasets to demonstrate the usefulness and promise of our proposed methods. Benefiting from IPS, increasing prompt point number results in significant performance gains for SAM, which may be further extended to include scrawl, sketch or coarse prompts. More importantly, our results prove that a fine-tuning-based approach is not necessarily optimal for utilizing a foundation model like SAM.
7 Limitations and future works
Although IPS has achieved promising performance in medical image segmentation tasks in endoscopic and dermoscopic modalities, the medical modalities involved in this paper still need to be increased to validate the effect of IPS in the entire medical image segmentation domain. Moreover, while our proposed APM approach can effectively improve the performance of transfer-free SAM in medical image segmentation, it performs sub-optimally when coupled with IPS. This suggests that we must trade between the end-to-end framework and performance.
Based on the above limitations, we plan to extend the ProMISe approach to as many medical image modalities ( including 2D and 3D) as possible. In order to improve the reliability of ProMISe, we plan to optimize the APM method while attempting a lightweight, non-invasive shifting of the representation encoder. Moreover, to improve the extensibility and practicality of ProMISe and IPS, we plan to enrich further the prompt types, such as bounding boxes, scribble, sketch, and text.